# Always Keep Your Promises: DynamicLRP, A Model-Agnostic Solution To Layer-Wise Relevance Propagation

### v1

@@ 4.3 The Promise System
When we reach a Node in traversal where propagation would halt from such a case, we instantiate a Promise, which defers the propagation computations and retrieves the missing tensors from further down in the graph.
Conceptually, a Promise acts as a placeholder for missing activations. When an operation requires an unavailable tensor to compute relevance, the Promise suspends its propagation but continues propagating through the graph to recover the needed values. Once these are found, the relevance propagation at the Node in question is computed, and the propagation "catches up" across all the Nodes that were traversed during the search, without backtracking and retraversal of the computation graph. We will now formally define this system.
@@ 4.3.1 Basic Promise Mechanism
Definition 2 (Promise).
A Promise is a mutable metadata object that is attached to a Node which requires some uncached forward pass input to compute its relevance propagation. The core structure of a Promise object is defined below:
      "rout": R_out_curnode,
      "args": [ None ] * num_missing_inputs,
      "rins": [ None ] * num_missing_inputs
The objective of a Promise is to act as a placeholder while we continue traversing the graph in search of themissing values. Promises have Promise Branches, which are where the metadata of the search and retrieval get stored, one for each missing input. A Promise’s Origin Node refers to the Node for which the Promise was created. A Promise Branch’s Arg Node refers to the first Node along a Promise Branch in which its forward pass output is retrievable from its own cached tensors.
Each intermittent Node between the Origin and Arg Nodes contributes two closures, one for its forward operation and one for its relevance rule. These form a pair of direction-based executable chains. When an Arg Node is reached, its forward output tensor is extracted and iteratively passed through the forward function chain to reconstruct the activation at the Origin Node, which is then stored within the Promise. Once all Arg Nodes of a Promise have been resolved, the Origin Node’s deferred relevance can be “fast-forwarded” down each Branch via its backward chain in the same manner. This mechanism effectively defers propagation until all dependencies are satisfied, ensuring continuity without backtracking or retraversal of the computation graph. We provide a full example of this mechanism in Appendix D.1.
@@ 4.3.2 Promise Nesting and Trees
When a Promise-Generating Node $v$ receives a Promise Branch $b_{P}$ as input, $v$ creates a new Promise $P^{\prime}$ and nests all of $P^{\prime}$’s branches as children of $b_{P}$ via parent-child connections, forming a Promise Tree. Promise Trees resolve bottom-up, starting with leaf Promise Branches, which encounter Arg Nodes, and forward-chain to reconstruct their ancestor Promises’ activations via parent connections. Once all activations of the Promise at the root of the Tree have been recovered, the relevance is propagated back down the Tree via child connections and backward chains. See Appendix D.2 for the formal resolution algorithm.
@@ 4.3.3 Promise Deadlock
Recall that our traversal heuristic states that we may only consider a node for traversal if all of its in-neighbours have propagated to it. Here, A creates a Promise $P$, propagates its Promise Branches $p_{1}$, $p_{2}$ to B and D, and we see that the number of inputs landed at B equals B’s in-degree. We traverse to B.
@@
B is an Arg Node, so the Promise Branch will retrieve its forward output and forward-chain it to obtain the activation at A, then save it to the Promise. We now stall traversal and enqueue B until the Promise is complete. Our only next option for traversal is D, which has received the other Promise Branch $p_{2}$ from A as input.
@@
We know that now E must be traversed, but see that we never traversed into it from the A->B->C path, so its incoming relevance inputs will never equal its in-degree, and we cannot consider it for traversal. This creates a circular dependency: C is stalled waiting for A’s Promise $P$ to complete, but $P$ cannot complete because its second branch $p_{2}$ requires accessing E’s activation, and E cannot be visited because it is waiting for C to propagate. The traversal is deadlocked.
The solution, in brief, is to break our traversal heuristic upon such situations, and allow Promise Branches to fully propagate until they reach Arg Nodes. The specific manner in which this is done is explained fully in Appendix D.3.
A simplified version of the propagation algorithm which considers all these is presented as Algorithm 2.
Note that this allows delayed computation to occur without the graph traversal pointer having to backtrack or revisit any node not on the current traversal frontier. Propagation across any given node only occurs once, either directly via tensor relevances or in a delayed manner via Promises. We claim that it follows that the internal nodes recorded within each Promise’s computation path are mutually exclusive among all Promises, though multple Promises may share the same Origin or Arg nodes. We prove this in Appendix E. This allows us to define clear upper bounds on the computational and memory overhead introduced by the Promise system.
4.4 Theoretical Analysis
We analyze the computational complexity of our promise-based LRP algorithm in terms of the architectural properties that determine promise resolution requirements.
Definition 4 (Promise Depth).
$d(v_{p})=\max_{u\in outadj(v_{p})}dist(u,argnode)$
Theorem 1 (Promise-Based LRP Complexity).
Let $\delta=\sum_{v_{p}\in V_{P}}d(v_{p})$
Proof.
@@ 4.4.1 Caching Promise Structures
This covers, however, only the first pass of the LRP algorithm. It is worth noting that repeated runs of LRP on the same model architecture yield an identical set of saved Promise computation paths. As a result, caching these paths can significantly improve the scalability of the algorithm. Once a Promise’s computation path has been defined, it can effectively be reduced to a single node, therefore eliminating the graph traversal overhead of its internal nodes.
Another benefit which arises from both the caching and this reduction is that the traversal order becomes strictly determined by the topological ordering of the nodes. Recall that due to the issue of Promise Deadlock that can be encountered during initial traversal, propagation of Promises may cause the traversal to temporarily disobey this ordering for the sake of searching ahead for missing activations. However, if we know the Promise chains beforehand and can reduce them to single nodes, then the order in which we perform computations becomes fully deterministic and independent of edge connections, reducing the graph traversal overhead to $O(n)$.
@@ 5 Experimental Setup
5.1 Models and Datasets
To evaluate faithfulness, we used finetuned versions of RoBERTa-large (335M) Liu et al. (2019) and Flan-T5-large (780M) Chung et al. (2022) with the SQuAD-v2 question-answering dataset Rajpurkar et al. (2018); Chan et al. (2021); Lee (2023) for text. For visual tests, we used the base ViT-b-16 Dosovitskiy et al. (2021) with the ImageNette-320 Howard (2019) test set and base VGG16 Simonyan & Zisserman (2015) from TorchVision on the CIFAR-10 ImageNet dataset Krizhevsky & Hinton (2009). The transformer models were all downloaded from HuggingFace. Finetuning settings can be found in respective model cards linked in citations.
5.2 Evaluating Attributions in Question-Answering Tasks
We measure the faithfulness of attributions via the accuracy of the top-1 and top-2 attributed tokens being within the answer span predicted by the model, and the accuracy of the top-1 token being within the span of the ground truth answer. We include also the Intersection-over-Union (IoU) between the top-2 attributed tokens and the model predicted span. The top-2 accuracy was chosen since QA models produce only two primary signals for the answer span, not necessarily one for all tokens in the span. We use these top-2 metrics only on with model predictions since LRP is meant to attribute the model’s decision process, not be a predictor of the answer itself. Though, alignment of the top-1 label accuracy and model exact match rate and F1 score can emphasize explainability strength of the attribution method. We make use of all answerable examples in the SQuADv2 validation set, and skip examples which the model blatantly predicts incorrectly (when the start index is after the end index). We run LRP-$\gamma$ with $\gamma_{linear}=0.001$ for both models, and only enable AttnLRP rules for the Flan-T5 attributions. See Table 1
5.3 Measuring Visual Attribution Faithfulness With Area Between Perturbation Curves
We measure faithfulness using the Area Between Perturbation Curves (ABPC) Schulz et al. (2020); Blücher et al. (2024). We first compute the relevance attributions for each input feature on each example. Then for $n$ iterations, we compute the Most-Relevant First (MoRF) curve at x = i by masking/occluding the top-i most relevant features from each input sample. We also do this inversely, masking/occluding the top-i least relevant features from each input sample, to compute the Least-Relevant First (LeRF) curve. The ABPC is then the difference of their integrals, or AUC(LeRF)AUC(MoRF)AUC(LeRF)-AUC(MoRF). All image inputs are 3-channel and 224x224 pixels in size. For each step, we occlude 1024 pixels (for the sake of computation time) from each image by replacing them with their values in a Gaussian-blurred version of the image ($\text{kernel}size=51$, $\sigma=20$). For ViT-b-16, we occlude the top 4 16x16 patches pooled by max relevance at each step, and for VGG16 we occlude the individual pixels. See Table 3, Table 3.
@@ 6 Results
The text experiments of Table 1 show DynamicLRP attributions to be strong explainers of model behaviour in LLM question answering. We compare these to the results obtained by Achtibat et al. from LxT (0.94 Top-1 accuracy, 0.84 IoU), though in computing IoU, they consider hits among the entire ground truth mask, whereas we compare against the model predicted mask and then only its endpoints. Meanwhile, we see that among all the attribution methods tested in Tables 3 and 3, ours and AttnLRP stand among the top of the ABPC metrics. Comprehensiveness is defined as AUC(y=baseline)AUC(MoRF)AUC(y=baseline)-AUC(MoRF) and Sufficiency as AUC(y=baseline)AUC(LeRF)AUC(y=baseline)-AUC(LeRF).
Table 1: DynamicLRP Attribution faithfulness measured by top attributed token accuracy against model predictions and labels and IoU with predictions with the entire SQuADv2 validation set (5928 examples). EM = Exact Match. Higher is better for all but last two columns. Columns marked with (P) are w.r.t. model predictions, (L) are w.r.t. labels. Top-2 accuracy and IoU values in parentheses signify strict matching to only the predicted start/end indices, and none of the indices between them. Non-parenthesized values here allow hits on indices within the span. All remaining unaccounted examples in the totals were skipped due to exceeding model max token length (512).
Model EM F1 top-1 (L) top-1 (P) top-2 (P) IoU (P) Examples Skipped
RoBERTa-L 82.29 86.27 93.70 97.47 96.22 (88.95) 92.70 (80.09) 5844 45
Flan-T5-L 83.65 90.15 95.06 98.64 94.73 (85.64) 89.94 (74.86) 5830 36
We also demonstrate that DynamicLRP is as faithful than the leading implementation of AttnLRP in the ViT experiment, and much better than all other alternative methods in every category present, barring the speed of InputXGradient. In addition, DynamicLRP outperforms the Zennit Anders et al. (2021) implementation, SmoothGrad, and Saliency in the VGG evaluations.
Table 2: ABPC and attribution efficiency metrics for ViT-b-16 experiments. 1000 examples from CIFAR-10 test set.
IG 0.10 0.95 0.85 N/A* >12GB
InputXGrad -0.0030 0.90 0.90 59.89 2.1GB
GradSHAP 0.074 0.94 0.87 3.32 8.6GB
Random -0.023 0.88 0.90 - -
AttnLRP 1.46 1.78 0.33 6.02 0.96GB
Ours 1.46 1.78 0.32 9.74 1.6GB
Table 3: ABPC and attribution efficiency metrics for VGG experiments. 3800 examples from ImageNette (320px) validation set.
Saliency 0.73 2.58 1.85 156.83 0.89GB
SmoothGrad 1.50 2.87 1.37 8.90 0.90GB
Random -0.0022 2.40 2.40 - -
Zennit LRP 1.69 2.72 1.02 29.41 1.4GB
Ours 1.77 2.65 0.88 21.79 2.3GB
What we also demonstrate is the framework’s capability to adopt rules defined by other implementations. In the ViT experiment, we tuned our LRP to behave similarly to specifications defined in AttnLRP Achtibat et al. (2024), using $\gamma$-LRP for Convolutional and Linear layers ($\gamma_{conv}=130$, $\gamma_{linear}=0.001$), while routing relevance only through the final matrix multiplication of Attention modules and skipping the softmax.
6.1 Architecture-Specific Factors on LRP Performance
Definition 5 (Promise Density).
The promise density of a neural network architecture is defined as:
$\rho=\frac{|V_{P}|}{|V|}$
where $|V_{P}|$ is the number of promise-generating operations and $|V|$ is the total number of operations in the computation graph.
We present in Table 4 the $d_{model}$, model depth, number of Promises, number of internal nodes in Promise paths, average Promise path length, Promise Density, and number of autograd nodes specific to each tested architecture.

### v3

@@ 4.2 The Promise System
When we reach a Node in traversal where propagation would halt from such a case, we instantiate a Promise, which defers the propagation computations and retrieves the missing tensors from further down in the graph. Conceptually, a Promise acts as a placeholder for missing activations. When an operation requires an unavailable tensor to compute relevance, the Promise suspends relevance propagation but continues traversing through the graph to recover the needed values. Once these are found, the relevance propagation at the problematic Node is computed, and the propagation “catches up” across all the Nodes that were traversed during the search, which are recorded to avoid backtracking and retraversal of the computation graph. We will now formally define this system.
@@ 4.2.1 Basic Promise Mechanism
Definition 1.
Let $\mathcal{T}$ be the space of tensors. We define a Promise $\mathcal{P}$ at a Node $u$ as a tuple $\mathcal{P}_{u}=(R_{out},\rho,\chi,\mathcal{A},\mathcal{R}_{in})$ where:
  • $R_{out}\in\mathcal{T}$ is the accumulated output relevance.
  • $\rho\in\{0,1\}$ is the readiness flag, where $\rho=1$ indicates that all missing arguments have been retrieved.
  • $\chi\in\{0,1\}$ is the completeness flag, where $\chi=1$ indicates that relevance has been distributed to all the inputs.
  • $\mathcal{A}=(a_{1},\dots,a_{k})$ is a sequence of retrieved argument tensors, where $a_{i}\in\mathcal{T}$.
  • $\mathcal{R}_{in}=(r_{1},\dots,r_{k})$ is a sequence of computed input relevances, where $r_{i}\in\mathcal{T}$.
The node $v$ where a Promise is instantiated is termed the Origin Node. To resolve the missing arguments $\mathcal{A}$, the Promise spawns $k$ Promise Branches, where the $i$-th branch is responsible for retrieving $a_{i}$. As an illustration, consider the following subgraph requiring a single Promise Branch, where red nodes do not store forward activations but require them for relevance propagation, green nodes do not store any forward activations and do not require them for relevance propagation, and yellow nodes store forward activations:
ABCDEModel forwardDo not store forward activations
Note that without node A’s forward activation, it is not possible to continue relevance propagation. However, since there exists a node, e.g. node $E$, further down the traversal path whose forward output is available, we can reapply the transformations of the nodes in between A and E to reconstruct A’s missing forward activation and propagate the relevance. For deferred propagation, each branch traverses the graph until it encounters a node $u$ whose forward output is locally retrievable; We term this the Arg Node.
Proposition 1: For any node $v$ in a computation graph, there exists at least one Arg Node which is a descendent of $v$ in the computation graph.
The proof is provided in Appendix 1.
Let $\pi=(v,w_{1},\dots,w_{m},u)$ be the traversal path from Origin $v$ to Arg Node $u$. For each intermediate node $w_{j}$, we record its forward operation $f_{w_{j}}$ and relevance rule $b_{w_{j}}$, forming two functional chains: a Forward Chain $\Phi_{v\leftarrow u}=f_{w_{1}}\circ\dots\circ f_{w_{m}}$ and a Backward Chain $\Psi_{u\to v}=b_{w_{m}}\circ\dots\circ b_{w_{1}}$. Upon reaching $u$, we retrieve its output $y_{u}$ and reconstruct the missing argument $a_{i}\in\mathcal{A}$ via $a_{i}=\Phi_{v\leftarrow u}(y_{u})$. Once all arguments in $\mathcal{A}$ are resolved, we set $\rho_{u}=1$, and the Origin Node $v$ computes the local relevance distribution $\mathcal{R}_{in}$ using $\mathcal{A}$ and $R_{out}$. Finally, the input relevance $r_{i}\in\mathcal{R}_{in}$ is fast-forwarded back to $u$ via the backward chain: $R_{u}=\Psi_{u\to v}(r_{i})$. This mechanism effectively defers propagation until dependencies are satisfied without backtracking. We provide a full example in Appendix C.1.
@@ 4.2.2 Promise Nesting and Trees
When a Promise-Generating Node $v$ receives a Promise Branch $p$ as input, $v$ creates a new Promise $\mathcal{P}^{\prime}$ and nests all of $\mathcal{P}^{\prime}$’s branches as children of $p$ via parent-child connections, forming a Promise Tree. Promise Trees resolve bottom-up, starting with leaf Promise Branches, which encounter Arg Nodes, and forward-chain to reconstruct their ancestor Promises’ activations via parent connections. Once all activations of the Promise at the root of the Tree have been recovered, the relevance is propagated back down the Tree via child connections and backward chains. See Appendix C.2 for the formal resolution algorithm.
@@ 4.2.3 Promise Deadlock
Recall that we may only consider a node for traversal if all of its in-neighbours have propagated to it. In Step (a), A creates a Promise $\mathcal{P}$, propagates its Promise Branches $p_{1}$, $p_{2}$ to B and D. We traverse to B, as all its inputs have now landed.
B is an Arg Node, so the Promise Branch will trigger a forward-chain to obtain the activation at A and save it. In Step (b), we now stall traversal and enqueue B until the Promise is complete. Next, in Step (c), we traverse to D, which received Promise Branch $p_{2}$ from A as input.
@@
Figure 2: Illustration of Promise Deadlock. (a) Origin node A distributes promise branches to retrieve activations. (b) Branch $p_{1}$ reaches Arg Node B, which stalls waiting for completion. (c) Branch $p_{2}$ reaches Arg Node D, which propagates to E. However, E depends on C, which depends on the stalled B, creating a cycle preventing Promise completion.
The solution, in brief, is to break our traversal heuristic upon such situations, and allow Promise Branches to propagate until they reach Arg Nodes. The specific manner in which this is done is explained fully in Appendix C.3, but a simplified version of the propagation algorithm that considers all of them is presented in Algorithm 2. Note that this allows delayed computation to occur without the graph traversal pointer having to backtrack or revisit any node not on the current traversal frontier, a property which we prove in Appendix E and allows us to define clear upper bounds on the computational and memory overhead introduced by the Promise system in Appendix 1.
Algorithm 2 Operation-Level Relevance Propagation With Promises
1:Input:Model output hidden_states, Relevance attribution target target_node, Incoming relevance $R_{out}$
2:Initialize $\texttt{stack}\leftarrow[\texttt{hidden\_states.grad\_fn}]$
3:Initialize $\texttt{heuristic\_break\_queue}\leftarrow[\hskip 1.84995pt]$
4:Initialize $\texttt{input\_tracker}\leftarrow\{\texttt{node}:[\hskip 1.84995pt]\text{ for all nodes in graph}\}$
5:Initialize $\texttt{nodes\_pending}\leftarrow\{\texttt{node}:\text{in-degree of }\texttt{node}\text{ for all nodes in graph}\}$
6:repeat
7:  if heuristic_break_queue is not empty then
8:    $\texttt{curnode}\leftarrow\texttt{heuristic\_break\_queue.dequeue()}$
9:    Create new Pre-Promise $P_{pre}$, assign Promise Branch in input_tracker[curnode] as parent $\triangleright$ See line 30
10:    Store $P_{pre}$ in curnode metadata
11:    $\texttt{curnode\_in\_rel}\leftarrow P_{pre}$
12:  else
13:    $\texttt{curnode}\leftarrow\texttt{stack.pop()}$ $\triangleright$ Process next node
14:    $\texttt{curnode\_in\_rel}\leftarrow\text{Relevance accumulated at }\texttt{curnode}$
15:    if curnode is target_node then $\triangleright$ Target reached
16:     return $R_{in}:=\texttt{curnode\_in\_rel}$
17:    end if
18:  end if
19:  if curnode requires Promise then $\triangleright$ Defer computation
20:    Create new Promise $P$ using type(curnode) and assign parent-child connections if applicable
21:    Assign each child of curnode its respective branch of $P$ via input_tracker[child]
22:  else
23:    Execute propagation function for curnode $\triangleright$ Applies LRP rule
24:    Assign each child of curnode its respective split of the relevance via input_tracker[child]
25:  end if
26:  for each child of curnode do
27:    Decrement nodes_pending[child] by 1 $\triangleright$ Update dependencies
28:    if nodes_pending[child] is 0 then
29:     Push child to stack $\triangleright$ Enqueue ready children
30:    else if input_tracker[child] contains a Promise Branch then
31:     Add child to heuristic_break_queue
32:    end if
33:  end for
34:until stack is empty $\triangleright$ All nodes processed
35:curnode $\leftarrow$ stack.pop() $\triangleright$ Process next node
36:return Input relevance $R_{in}$
@@ 4.2.4 Caching Promise Structures
The previously described methodology covers only the first pass of the LRP algorithm. Repeated runs of LRP on the same model architecture yield an identical set of saved Promise computation paths. So, we cache these paths to improve the algorithm’s scalability. Once a Promise’s computation path has been defined, it can effectively be reduced to a single node, eliminating the graph traversal overhead of its internal nodes. Finally, consider that if we know the Promise chains beforehand and can reduce them to single nodes, then the order in which we perform computations becomes fully deterministic and independent of edge connections, reducing the graph traversal overhead to $O(n)$, even when promise deadlock is encountered during initial traversal.
@@ 5 Experimental Setup
In this work, we aim to demonstrate that DynamicLRP produces faithful explanations for modern architectures of varying scale. In attempt to measure faithfulness, we adapt the evaluation strategies from Achtibat et al. (2024) using classification and exctractive question-answering tasks.
5.1 Classification Tasks
Input perturbation, as explored and advanced by Blücher et al. (2024), DeYoung et al. (2020), and Schulz et al. (2020), has become a widely-used standard for measuring attribution faithfulness. This method leverages the model itself as an oracle to validate the feature ranking generated by an attribution method. By systematically removing or replacing input features (e.g., tokens or pixels) and observing the degradation in performance with respect to a baseline, we can verify if a feature’s attributed “importance” is in-line with a model’s prediction. We quantify this using the Area Between Perturbation Curves (ABPC), ranking features by importance and iterating in both orders:
  1. 1.
    Most Relevant First (MoRF): We occlude the most important features first. A faithful attribution should result in a steep decline in the model’s confidence for the true label, as crucial information is removed.
  2. 2.
    Least Relevant First (LeRF): We occlude the least important features first. A faithful attribution should result in a flat curve or minimal decrease (potentially even an increase) in confidence, as irrelevant features are removed.
The ABPC is defined as $\text{AUC}(\text{LeRF})-\text{AUC}(\text{MoRF})$, forming a unified metric where higher values indicate better faithfulness (higher stability under LeRF and sharper drop under MoRF). Other metrics related to faithfulness are comprehensiveness, defined as $\text{AUC}(\text{baseline})-\text{AUC}(\text{MoRF})$ and sufficiency, defined as $\text{AUC}(\text{baseline})-\text{AUC}(\text{LeRF})$.
Visual Classification Setup
We use the the base VGG16 model (Simonyan and Zisserman, 2015) evaluated on the ImageNette-320 dataset (Howard, 2019), and the base ViT-b-16 (Dosovitskiy et al., 2021) architecture on the CIFAR-10 dataset (Krizhevsky and Hinton, 2009). For perturbation, we occlude features by replacing them with a Gaussian-blurred version of the image ($\text{kernel size}=51$, $\sigma=20$). For ViT-b-16, we occlude the top-4 $16\times 16$ patches (pooled by max relevance) at each step. For VGG16, we occlude 1024 individual pixels per step.
Text Classification Setup
Following the experimental setup in Achtibat et al. (2024), we first use a finetuned Llama-3.2-1B model (Grattafiori et al., 2024) on the IMDb dataset (Maas et al., 2011) for sentiment analysis. We evaluate explanations by determining the relevance of input tokens to the predicted class. For language modelling, we use the base model on the English subset of the Wikipedia  [^1] Foundation dataset and frame the task as next-token prediction. We compute attribution for the ground-truth next token given a context window of 512 tokens. In both settings, we apply LRP-$\gamma$ with $\gamma_{linear}=1.0$ and compare against baseline attribution methods using MoRF and LeRF occlusion metrics after masking one token per step.
5.2 Extractive Question-Answering Tasks
Extractive Question-Answering (EQA) tasks differ structurally from classification: the model identifies a span within the input context as the answer. In this setting, the MoRF/LeRF strategy can be redundant, as the tokens constituting the predicted span are by definition necessary for the model to output that specific span. Therefore, we simplify the measurement to checking if the highest-ranked tokens are within the predicted answer span. We evaluate faithfulness via the accuracy of the top-1 attributed token falling within the label golden span and the model’s predicted span. We use finetuned versions of RoBERTa-large (Liu et al., 2019) and Flan-T5-large (Chung et al., 2022) on the SQuAD-v2 dataset (Rajpurkar et al., 2018; Chan et al., 2021; Lee, 2023). We include all answerable examples in the validation set, skipping only those where the model predicts an invalid span (start index > end index). We run LRP-$\gamma$ with $\gamma_{linear}=0.001$. We mask the relevance of tokens found in the question before taking the top-1 attributed token.
@@ 6 Results
As can be seen from tables 1, A.2, and A.1, DynamicLRP consistently outperforms generic attribution methods in vision tasks and matches or surpasses model-specific AttnLRP implementation (Figure 3).
Figure 3: Visual representation of attributions produced by different LRP algorithms for the predicted class truck by a ViT-b architecture. (a) Original image (b) Zennit LRP attributions (c) DynamicLRP attributions
In Language tasks, DynamicLRP demonstrates faithful attribution in all cases, achieving the best score in three out of five language tasks and ranking second in the others. When compared on the larger model, LLaMA3.2-1B, we note that the specialized Attn-LRP implementation outperforms DynamicLRP. We attribute this difference to relevance distribution rather than signal quality. DynamicLRP tends to spread relevance more broadly across influential euqally influential tokens, whereas AttnLRP produces sharper peaks that are favoured by rank-based metrics. Qualitatively, however, both methods yield highly consistent attribution signals and identify the same semantic features as important (Figure A.2). For extractive question answering tasks, DynamicLRP outperforms all other methodologies. It is worth mentioning that Top-1 Token-to-Gold-Span(TGS) accuracy is tightly coupled to model predictive performance and therefore confounds attribution quality with task accuracy, requiring careful interpretation alongside metrics such as F1 and Exact Match (Table A.3). To disentangle these effects, we consider Top-1 Token-to-Predicted-Span (TPS), which measures whether attributions identify the tokens actually used by the model for its prediction. LRP-based methods maintain consistently high TPS, while gradient-based methods degrade sharply on erroneous predictions. This indicates that LRP preserves the importance of the predicted span throughout the computation path, even when the model fails. The robustness of DynamicLRP on variuos models highlights its practical value as an architecture-agnostic diagnostic tool, particularly during model development, where explaining failures is often more informative than explaining successes.
Table 1: Quality of explanations across attribution methods. For vision and language classification tasks(ImageNette, CIFAR, IMDb and Wikipedia), faithfulness is measured by the Area Between MoRF and LeRF perturbation curves. For question answering (SQuADv2) tasks, we report Top-1 Token-to-Golden-Span (TGS) and Token-to-Predicted-Span (TPS). Higher is better; best results in bold.
Task type Image Class. Text Class. Causal LM Extractive QA
Method VGG ViT-b-16 LlaMa3.2-1B LlaMa3.2-1B BERT RoBERTa-L T5-Flan-L
ImageNette-320 CIFAR-10 IMDB Wikipedia SQuADv2 SQuADv2 SQuADv2
IG 1.39 0.10 1.77 1.68 0.438 (0.558) 0.353 (0.134) 0.798 (0.127)
GradSHAP 1.62 0.074 2.61 1.59 0.419 (0.527) 0.374 (0.143) 0.795 (0.157)
Zennit (Attn-LRP) 1.69 1.46 3.73 3.16 0.517 (0.805) N/A N/A
DynamicLRP (ours) 1.77 1.46 3.18 2.59 0.579 (0.986) 0.937 (0.889) 0.951 (0.986)
The architecture is not supported by the respective method.
6.1 Architecture-Specific Factors on LRP Computational Overhead
DynamicLRP’s computational overhead depends on two key architectural traits: Promise Depth ($\delta$) and Promise Density ($\rho$). Promise depth $\delta$ of an architecture is the maximum graph distance to an Arg Node for all Promise Branches. The Promise density is the ratio of promise-generating operations to total operations, $\rho=\frac{|V_{P}|}{|V|}$. Table 2 reports these values across architectures, providing empirical insight to the runtime and memory complexity. Notably, model variants of Flan-T5-large show that increased model complexity (e.g., more autograd nodes) does not necessarily scale up Promise density. This suggests that complex training procedures or architectural expansions do not disproportionately increase the “promise overhead” implying favourable scalability for our method.